OpenStudy (anonymous):

Decide whether Y is a function of X y=-2x-4

OpenStudy (shayaan_mustafa):

yes. because as you changes x then y changes. it implies that y depends on x. so y is the function of x. for example. substitute x=1 y=-2-4 y=-6 now substitute x=2 y=-4-4 y=-8 see. as x is changes then so y. I hope this help you.